What is the value of x in the equation 4x+8y=40 when y=0.8
Answer by xxLaurax(5)

So, you are trying to get x by itself. So, the first thing you do is plug in 0.8 for y into the equation.
4x+8(0.8) = 40
4x+6.4 = 40 Then you subtract 6.4 from both sides.
40-6.4 = 33.6
4x=33.6 Now get the x by itself by dividing 4 on both sides.
33.6/4 = 8.4
x = 8.4 is your answer.